f (x)= x +5 and g (x) =|x |, find (f• g)(-4) and (g• f)(-4)
Answer by Tatiana_Stebko(1539)

(f• g)(x)=f(g(x))=g(x) +5=|x |+5
(f• g)(-4)=|-4|+5=4+5=9
(g• f)(x)=g(f(x))=|f(x) |=|x+5|
(g• f)(-4)=|-4+5|=|1|=1